OpenStudy (anonymous):

Kendra takes bread to her two friends' houses every week. She takes 3 times the amount to Judy as she does Lynn. Combined, she takes them 12 loafs. How much does she give Judy?

OpenStudy (anonymous):

3x + x = 12 4x = 12 x = 3

OpenStudy (anonymous):

If she takes 3 times the amount to Judy, she takes 3x, and we can't forget that this is a total of both breads, so + x would be the bread she gives to Lynn.

OpenStudy (anonymous):

You can plug it into the equation to check if it's correct. 3(3) + 3 = 12 9 + 3 = 12 12 = 12 TRUE

OpenStudy (triciaal):

@Juanita How much does she give Judy? is the original question so be careful not to switch